OpenStudy (anonymous):

I need help with this too.. sketch the graph for f(x ) =x2-8x+16?

OpenStudy (owlcoffee):

are you familiar with differential calculus?

OpenStudy (anonymous):

this is a parabola to start off with

OpenStudy (zale101):

f(x ) =x2-8x+16 You can either plot points, or convert the equation into vertex form to and find the y intercepts of the function. Vertex form: y=(x-h)^2+k Vertex point is (h,k) The first step to convert a strandard quadratic equation into vertex form is by completeing the square. Are you familiar with that?

OpenStudy (anonymous):

it is an upward parabola,complete the square to find vertex

OpenStudy (zale101):

Yes, all quadratic equations are parabolas in graphs.

OpenStudy (anonymous):

no wrote it out please?

OpenStudy (anonymous):

work

OpenStudy (zale101):

First, do you know what is the vertex point on the parabola?

OpenStudy (zale101):

Or, how to find the x and y intercept of the function?

OpenStudy (owlcoffee):

@Zale101 are you sure he's have to apply a parabola equation from analytic geometry? I believe it's unnecesaryc complex.

OpenStudy (anonymous):

no

OpenStudy (zale101):

How to complete the square f(x) =x^2-8x+16 \[y=x^2-8x+16\] Step 1: subtract 16 to both sides \[y-16=(x^2-8x)\] Step 2:Take the half of the coefficent, which is 8, and then square it

OpenStudy (zale101):

\((\frac{ 8 }{ 2 })^{2}\)=?

OpenStudy (anonymous):

8

OpenStudy (zale101):

(8/2) is 4, you square it 4^2=16 :) so 8 was the wrong answer?

OpenStudy (zale101):

@owlcoffee i believe this question is algebra based. @keyshia.hart1 is that right?

OpenStudy (zale101):

Anyways, back to your problem. So, we got for (8/2)=4 and 4^2=16, you agree @keyshia.hart1 ?

OpenStudy (zale101):

\[(x^2-8x+16)=y-16-16\] Now, factor (\(x^2-8x+16)\) what would you get?

OpenStudy (zale101):

http://www.purplemath.com/modules/sqrquad.htm If you have no idea what i'm talking about, maybe this would help

OpenStudy (anonymous):

im lost

OpenStudy (owlcoffee):

In order to graph a function, we will have to first study the domain of the function, in other words, the values of x that can cause an indetermination on the function. Since it's a quadratic function, we can be sure that no value of x will cause an indetermination on the domain, we write it like this: \[d(f)= \mathbb{R} \] Second, let's analyze the "roots" of the function, and later on, the y-interception. Let's first begin with the roots of the function, we will call "root" to all those values of x that make the function equal zero, in other words, the x-axis interception. so: \[x^2 -8x+16=0\] this is a quadratic equation, so we'll use the general formula to find the x values that satify it: \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \[x=\frac{ 8 \pm \sqrt{8^{2}-4(1)(16)} }{ 2(1) }\] \[x=\frac{ 8 \pm \sqrt{0} }{ 2 }\] \[x=4\] since we got a square root of zero, means that the function will be tangent to the x-axis. Now, since we got the roots, we can study the sign of the function, with a line that represents the values of the image of the x axis. Since the squared component of the function is positive, I can be certain that the images to the right (x-values go bigger) the value of the function will remain positive:  it's positive on both sides, because we had a double root. A double root is a special case, and indicates a point of tangency to the x-axis. So, we can discard that the function would ever take negative values. let's now find the "y-interception". The y interception is the image of the function when x=0. therefore: \[f(0)=(0)^2-8(0)+16\] \[f(0)=16\] now, with all the information found, we can now sketch it:  (just imagine that it's well drawn)